OpenStudy (anonymous):

I need help rewriting y=x^2+14x+50 in vertex form .

OpenStudy (ranga):

Complete the square.

OpenStudy (ranga):

y=x^2+14x+50 Take the coefficient of x which is 14 and divide it by 2. 14/2 = 7 This 7 will go inside the square. Then you subtract 7^2: y = (x + 7)^2 - 7^2 + 50 = (x + 7)^2 - 49 + 50 y = (x + 7)^2 + 1 That is the vertex form.

OpenStudy (anonymous):

Thank You So Much @ranga !!!!! Your great

OpenStudy (ranga):

You are welcome.